OpenStudy (anonymous):

Points (5,3) and (6,2) lie on line k. Which is the slope of the line that is parallel to k?

OpenStudy (anonymous):

the slope will be the same on the original slope

OpenStudy (anonymous):

like\[2 - 3 / 6 - 5 = -1/ 1 = -1\] so the parallel line's slope will be -1

OpenStudy (anonymous):

thank you friedrice :)

OpenStudy (anonymous):

your welcome